OpenStudy (anonymous):

find an equation of the line having the given slope and containing the given point m=7/8,(6,-8)

OpenStudy (btaylor):

in slope-intercept form (y=mx+b), plug in your known values of y, x, and m. Solve for b. YOu get the equation.

OpenStudy (anonymous):

Use the form y = mx + b m = slope b = y-intercept Plug in what you have. y = 7/8x + b Plug in the coordinate on the line. -8 = 7/8(6) + b -8 = 21/4 + b -53/4 = b y = 7/8x - 53/4

OpenStudy (anonymous):

can you simplify your answer

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

my fault numbers should have read m=3/4x,(6,-8)

OpenStudy (anonymous):

m=3/4,(6,-8)

OpenStudy (anonymous):

Do the same thing. y = 3/4x + b -8 = 3/4(6) + b -8 = 9/2 + b -25/2 = b y = 3/4x - 25/2

OpenStudy (anonymous):

thanks calcmathlete can you help me on all my math problems

OpenStudy (anonymous):

are you free for the rest of the the night

OpenStudy (anonymous):

Sorry. I have to go in a few minutes.

OpenStudy (anonymous):

what time will you be back on here

OpenStudy (anonymous):

Not sure...

OpenStudy (anonymous):

alright thanks for your help

OpenStudy (anonymous):

You're welcome